The degree of operating leverage for Williams Company is 5. The actual operating income is $20,000. If the company expects a 45% increase in sales, operating income should increase by $9,000 O True O False
Answer #1
The statement is false which is shown as below:-
Operating leverage = (%change in operating income/ % change in sales)
= 5 = %change in operating income / 45%
= %change in operating income = (45*5) = 225%
Hence, change in operating income is (20,000*225%) = $45,000